Circular No. 6323

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

NOVALIKE VARIABLE IN SAGITTARIUS S. Nakano, Sumoto, Japan, forwards the following precise position from

Y. Kushida, Yatsugatake South Base Observatory, obtained on Feb. 23.792 UT with a 0.25-m reflector (+ CCD): α = 17^{h}52^{m}32^{s}.69, δ = −17^{o}41′07′′.7 (equinox 2000.0).

GRO J1744–28 D. A. Frail, Very Large Array (VLA), National Radio Astronomy Ob-

servatory; C. Kouveliotou, Universities Space Research Association; J. van Paradijs, University of Alabama at Huntsville and University of Amsterdam; and R. Rutledge, Department of Physics, Massachusetts Institute of Technology, report: “We have continued to monitor the time-variable radio source in the XTE error box of GRO J1744–28 (cf. IAUC 6309). Observations with the VLA at 3.6 cm were made on Feb. 12, 13, 16, and 17 for 5, 60, 22, and 15 min, respectively. Flux densities for these four days are < 345 ± 115, < 195 ± 65, < 165 ± 55, and 300 ± 55 μJy, respectively. It is clear from these and our earlier measurements (IAUC 6307) that the variations in the radio flux density are not due to a gradual rise or decay of a synchrotron source, but they appear to be stochastic. A weighted average of all three detections results in a location for the time-variable radio source of α = 17^{h}44^{m}36^{s}.820 ± 0^{s}.008, δ = −28^{o}45′37′′.38 ± 0′′.17 (equinox 2000.0), in close agreement with the position given on IAUC 6307. The radio source is offset 0′′.6 from the optical candidate identified on an ESO R plate (IAUC 6314) and 1′′.7 from star C (and star a) identified by Vanden Berk et al. (IAUC 6315) and Miller (IAUC 6318). We estimate that systematic errors (if any) in the VLA measurements can account for no more than 0′′.15 of this discrepancy.”

COMET C/1996 B2 (HYAKUTAKE) Total visual magnitude estimates (cf. IAUC 6315): Feb. 16.69 UT, 7.6

(T. Lovejoy, Jimboomba, Queensland, 15×80 binoculars); 18.21, 7.9 (A. Pereira, Cabo da Roca, Portugal, 9×34 binoculars); 20.29, 7.4 (J. G. de S. Aguiar, Campinas, São Paulo, Brazil, 11×80 binoculars); 21.30, 7.1 (G. W. Kronk, Troy, IL, 20×80 binoculars); 22.24, 7.4 (Pereira; 1^{o} tail in p.a. 320^{o}); 23.44, 7.4 (A. Hale, Cloudcroft, NM, 10×50 binoculars).

1996 February 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green